OpenStudy (anonymous):

3 hours and ten minutes to seconds. How to? Do i use 3.10 for the multiplication?

OpenStudy (anonymous):

An hour is \(60\times60=2600\) seconds A minute is \(60\) seconds.

OpenStudy (anonymous):

You could convert the minutes into hours, giving you: \[ 3+\frac{10}{60} \]

OpenStudy (anonymous):

But \(3.10\) is wrong because \(10/60\neq 10/100 = 0.10\)

OpenStudy (anonymous):

Oh. i see! So when ive calculated the hours. then for the 10 minutes, i just multiply 60 x 10

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Oh ok Thank you!

OpenStudy (anonymous):

Also I should have said \(60\times 60=3600\)

OpenStudy (anonymous):

Oh great thanks!